OpenStudy (anonymous):

What is the volume of the cube? A. 6 cubic inches B. 8 cubic inches C. 12 cubic inches D. 16 cubic inches

OpenStudy (anonymous):



OpenStudy (anonymous):

@jjuden

OpenStudy (anonymous):

sorry the drawing sucks

OpenStudy (jjuden):

V=a^3

OpenStudy (anonymous):

so V=..16?

OpenStudy (anonymous):

no its 8

OpenStudy (jjuden):

2*2=4 4*2=8 8 is your answer

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

A cube has a side length of 6 feet. What is the volume of the cube? A. 6 ft3 B. 18 ft3 C. 36 ft3 D. 216 ft3 <-----i picked this one, is it right?

OpenStudy (anonymous):

you would (ahem) CUBE it hence the name

OpenStudy (anonymous):

Cube it?